Hey, I've been set a chemistry test to finish for Monday and I'm struggling on these 2 questions.. any help would be great!

The first one.. what is the mass of Fe produced when 500g of iron oxide is used?

2Fe2O3 . → . 4Fe . + . 3O2

Calculate molar mass of Fe2O3
Calculate the number of moles of Fe2O3
What is the molar ratio?
Calculate the number of Fe moles
What is the mass of Fe produced?

And the second question... Calculate the mass of NaOH required to produce 20g of Na2SO4

2NaOH + H2SO4 → Na2SO4 + 2H2O

If you could tell me the equations to reach the answers that'd be great, thank you!
1
Expert's answer
2017-01-23T06:45:49-0500
Solution.
a) 2Fe2O3 → 4Fe + 3O2
M(Fe2O3) = 562 + 163 = 160 g/mol
n(Fe2O3) = m/M = 500/160 = 3.125 mol
(2160) g Fe2O3 ‒ (456) g Fe
500 g Fe2O3 ‒ x g Fe
x = 350 g Fe
n(Fe) = m/M = 350/56 = 6.25 mol
Molar ratio Fe2O3 to Fe is 3.125/6.25 = 1/2

b) 2NaOH + H2SO4 → Na2SO4 + 2H2O
M(NaOH) = 23+16+1 = 40 g/mol
M(Na2SO4) = 232 + 32 + 164 = 142 g
(240) g NaOH ‒ (142) g Na2SO4
x g NaOH ‒ 20 g Na2SO4
x = 11.27 g Na2SO4

Answer: a) m(Fe) = 350 g
b) m(Na2SO4) = 11.27 g